Solve
2x^2-y^2=1
x+y=12
Please explain how to solve this in any form.
Answer by mananth(16946) 
 2x^2-y^2=1
x+y=12
x=12-y
2x^2-y^2=1
substitute x
2(12-y)^2-y^2=1
2(144-24y+y^2)-y^2=1
288-48y+2y^2-y^2=1
y^2-48y+288=1
y^2-48y+287=1
y^2-41y-7y+287=0
y(y-41)-7(y-41)=0
(y-41)(y-7)=0
y=41 OR 7
When y =41 x=-29 ( -29,41)
when y=7, x=5 (5,7)
|
|
|
| |